Grade 10General Physics

the linear momentum of a body P varies with time and is given by the eq. P=x +yt​2 where x and y are constants. the net force actiong on the body for one dimensional motion is proportion to:- (1) t (2) a constant (3) 1/t (4) none of these

Ishita Hariramani
10 Years agoGrade 10

1 Answer

Vikas TU
10 Years ago
hence first option.the linear momentum of a body P varies with time and is given by the eq. P=x +yt​2

Write down the change in momentum eqn.
.i.e differentiate P w.r.t  to time --------->
dP/dt = 0 + 2yt = Force which is directly proportional to t